If one side of a square is tripled and another is increased by 2, the area of the resulting rectangle is 45 units. What is the length of a side of the square?
Answer by JimboP1977(311)

3x*(x+2)=45
3x^2+6x=45
x^2+2x=45/3
(x+1)^2-1=45/3
x=sqrt(45/3)+1)-1
x=sqrt(16)-1
x=3